# On palindromic numerators of bigraded symmetric orbifold Hilbert series and Kostka-Foulkes polynomials

### v2

@@ 3.2 Harmonic polynomials of the symmetric group
The transformed Hall-Littlewood polynomials $H_{\lambda}(X;t)$ can also be expressed as graded characteristics of the representations of the symmetric group in certain spaces of polynomials [19]. Recall that as a bijection between symmetric functions and characters of representations of the symmetric group, the Frobenius characteristic is the linear map from the representation ring of characters of symmetric groups $R(S_{n})$ to the ring of symmetric functions which sends the irreducible representation indexed by a partition $\lambda$ to the Schur function $s_{\lambda}$. For any $S_{n}$-module $V$, denoting $\chi$ the character of a representation $V$, the Frobenius characteristic map of $V$ is
@@
$\mathcal{F}_{t}(V)=h_{n}\left(\frac{X}{1-t}\right).$ (33)
@@
is satisfied. The link between the companion ring $\mathcal{H}_{S_{n}}$ to the ring of $S_{n}$-invariants is established through a result of Chevalley (valid for any reflection group) [20], which states that there is a natural isomorphism of graded $S_{n}$-modules
@@
$\mathcal{F}_{t}\left(\mathcal{H}_{S_{n}}\right)=h_{n}\left(\frac{X}{1-t}\right)\prod_{k=1}^{n}\left(1-t^{k}\right)=\sum_{\left|\lambda\right|=n}K_{\lambda\left(1^{n}\right)}(t)s_{\lambda}\left(X\right),$ (38)
@@ 3.3 Palindromic numerators as a sum of products of Kostka-Foulkes polynomials
we are now ready to show that the palindromic numerators of the generating function $Z_{log}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q^{2}};q,\bar{q};\mathbb{C}^{2}\right)$ of bigraded symmetric orbifold Hilbert series can be expressed as sums of products of Kostka-Foulkes polynomials associated with a pair of partition $\lambda$ and $\mu=(1^{n})$, i.e, by relabeling the numerators as $\mathcal{P}_{n}\left(q,\bar{q}\right)$ that
@@
$h_{n}\left(\frac{X}{1-q}\right)\prod_{k=1}^{n}\left(1-q^{k}\right)=\sum_{\left|\lambda\right|=n}K_{\lambda\left(1^{n}\right)}(q)s_{\lambda}\left(X\right),$ (42)
according to Eq. (38), and using the Schur polynomial specialization [19]
@@
Using the hook-formula [8]
@@
which is in agreement with the numerators calculated in [5].
@@
In Eq. (50), we have used the well-known fact that the cycle index $Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)$ of the symmetric group $S_{n}$ in Eq. (49) can be stated in terms of Bell polynomials, which can be expressed as determinants.
@@ 4 Recurrences relations and differential operators
$\boxed{\mathcal{G}_{n+1}\left(q,\bar{q}\right)\cdot\mathcal{P}_{n+1}\left(q,\bar{q}\right)\cdot Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)=\mathcal{P}_{n}\left(q,\bar{q}\right)\cdot Z_{n+1}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n+1}\right).}$ (52)
From the recurrence relation (52), we write
@@
introduced in [1] as a multiplication operator whose action on $Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)$ gives
${\rm{\hat{X}}}Z_{n}=(n+1)Z_{n+1},$ (56)
@@
$\boxed{\mathcal{P}_{n+1}\left(q,\bar{q}\right)\cdot Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)=\left[\frac{\mathcal{P}_{n}\left(q,\bar{q}\right)}{(n+1)\mathcal{G}_{n+1}}{\rm{\hat{X}}}\right]\cdot Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)},$ (57)
@@
$\mathcal{D}=\frac{\mathcal{P}_{n}\left(q,\bar{q}\right)}{(n+1)\mathcal{G}_{n+1}}{\rm{\hat{X}}}$, (58)
@@ Acknowledgements
We thank the anonymous referee of an earlier version for encouraging us to move in the direction of the paper hereby presented. This work is supported by the South African Research Chairs initiative of the Department of Science and Technology and the National Research Foundation, and by the National Institute for Theoretical and Computational Sciences, NRF Grant Number 65212.

### v3

@@ 3.2 Harmonic polynomials of the symmetric group
The transformed Hall-Littlewood polynomials $H_{\lambda}(X;t)$ can also be expressed as graded characteristics of the representations of the symmetric group in certain spaces of polynomials [desarmenien1994hall]. Recall that as a bijection between symmetric functions and characters of representations of the symmetric group, the Frobenius characteristic is the linear map from the representation ring of characters of symmetric groups $R(S_{n})$ to the ring of symmetric functions which sends the irreducible representation indexed by a partition $\lambda$ to the Schur function $s_{\lambda}$. For any $S_{n}$-module $V$, denoting $\chi$ the character of a representation $V$, the Frobenius characteristic map of $V$ is
@@
$\mathcal{F}_{t}(V)=h_{n}\left[\frac{X}{1-t}\right].$ (33)
@@
is satisfied. The link between the companion ring $\mathcal{H}_{S_{n}}$ to the ring of $S_{n}$-invariants is established through a result of Chevalley (valid for any reflection group) [bergeron2009algebraic], which states that there is a natural isomorphism of graded $S_{n}$-modules
@@
$\mathcal{F}_{t}\left(\mathcal{H}_{S_{n}}\right)=h_{n}\left[\frac{X}{1-t}\right]\prod_{k=1}^{n}\left(1-t^{k}\right)=\sum_{\left|\lambda\right|=n}K_{\lambda\left(1^{n}\right)}(t)s_{\lambda}\left(X\right),$ (38)
@@ 3.3 Palindromic numerators as a sum of products of Kostka-Foulkes polynomials
We are now ready to show that the palindromic numerators of the generating function $Z_{log}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q^{2}};q,\bar{q};\mathbb{C}^{2}\right)$ of bigraded symmetric orbifold Hilbert series can be expressed as sums of products of Kostka-Foulkes polynomials associated with a pair of partition $\lambda$ and $\mu=(1^{n})$. By relabeling the numerators as $\mathcal{P}_{n}\left(q,\bar{q}\right)$, we claim that
@@
$h_{n}\left[\frac{X}{1-q}\right]\prod_{k=1}^{n}\left(1-q^{k}\right)=\sum_{\left|\lambda\right|=n}K_{\lambda\left(1^{n}\right)}(q)s_{\lambda}\left(X\right),$ (42)
according to Eq. (38), and using the Schur polynomial specialization [desarmenien1994hall]
@@
Using the hook-formula [macdonald1998symmetric]
@@
which is in agreement with the numerators calculated in [Mvondo-She:2019vbx].
@@
In Eq. (50), we have used the well-known fact that the cycle index $Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)$ of the symmetric group $S_{n}$ in Eq. (49) can be stated in terms of Bell polynomials, which can be expressed as determinants. We note that the determinant in Eq. (50) is essentially the determinant for $h_{n}$ in terms of power-sums (see page 28 in [macdonald1998symmetric]).
@@ 4 Recurrences relations and differential operators
$\mathcal{G}_{n+1}\left(q,\bar{q}\right)\cdot\mathcal{P}_{n+1}\left(q,\bar{q}\right)\cdot Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)=\mathcal{P}_{n}\left(q,\bar{q}\right)\cdot Z_{n+1}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n+1}\right).$ (52)
Then, from the recurrence relation (52), we write
@@
introduced in [Mvondo-She:2018htn] (also utilized in [Mvondo-She:2023ppz] in the context of urn models) as a multiplication operator [^2] We note that the operator $\hat{\mathrm{X}}$ when acting on power sums is the special case $\alpha=1$ of the operator on Jack polynomials found in Example 6, on page 385 of [macdonald1998symmetric]. which together with the lowering operator $\hat{D}=\frac{\partial}{\partial\mathcal{G}_{1}}$ generates the Heisenberg–Weyl algebra $\left[\hat{D},\hat{X}\right]=1$, and whose action raises the partition function $Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)$ to $Z_{n+1}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n+1}\right)$ as
${\rm{\hat{X}}}Z_{n}=Z_{n+1},$ (56)
@@
$\boxed{\mathcal{P}_{n+1}\left(q,\bar{q}\right)\cdot Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)=\left[\frac{\mathcal{P}_{n}\left(q,\bar{q}\right)}{\mathcal{G}_{n+1}}{\rm{\hat{X}}}\right]\cdot Z_{n}\left(\mathcal{G}_{1},\ldots,\mathcal{G}_{n}\right)},$ (57)
@@
$\mathcal{D}=\frac{\mathcal{P}_{n}\left(q,\bar{q}\right)}{\mathcal{G}_{n+1}}{\rm{\hat{X}}}$, (58)
@@ 5 Summary and outlook
We believe that from the results obtained in this article, an interesting story could emerge in our program of developing phenomenological applications of log gravity and ultimately understanding the AdS3/LCFT2 correspondence. Indeed, some light has been shed on the nature of what could be considered as $n$-particle highest weight states in the log sector, by proving their realization as (products of) harmonic polynomials. Furthermore, it appeared from [Mvondo-She:2024iop] that the log sector shows a remarkable dynamical self-organization into a solitonic configuration that forms a KP hierarchy. The sandpile model being a paradigmatic example of self-organized criticality, we proposed to study an analogy between sandpiles, which for various boundary conditions correspond to distinct LCFT descriptions, and log gravity. Harmonic polynomials have been shown to induce and to provide a framework to analyze sandpile dynamics [lang2019harmonic], and the results of this paper warrant further study in describing the algebraic structure of the log sector as a self-organized dynamical system.
@@ Acknowledgements
We thank the anonymous referees for various inputs that contributed to improving the presentation of this paper. This work is supported by the South African Research Chairs initiative of the Department of Science and Technology and the National Research Foundation, and by the National Institute for Theoretical and Computational Sciences, NRF Grant Number 65212.